# Global solutions to cross-diffusion systems with independent advections in one dimension

### 20 Mar 2026 v1

@@ 4.3. Existence of solutions in terms of $\pi_{t,x}$
Let $\pi_{t,x}$ be the Young measure of the sequence (4.1). Let $u$, $s$, $\overline{\frac{u}{s}\,\partial_{x}s^{\alpha}}$ and $\overline{\frac{v}{s}\,\partial_{x}s^{\alpha}}$ be limits as in (C1), (C2) and (C14). Then, for a.e. $(t,x)$
@@
In particular, suppose that for a.e. $(t,x)$
@@
This proves (4.15). Next, if (4.16) is satisfied then by (4.15), $\overline{\frac{u}{s}\,\partial_{x}s^{\alpha}}=\frac{u}{s}\,\partial_{x}s^{\alpha}$. By (C14), $\overline{\frac{v}{s}\,\partial_{x}s^{\alpha}}=\partial_{x}s^{\alpha}-\overline{\frac{u}{s}\,\partial_{x}s^{\alpha}}$ and we also have $\frac{v}{s}\,\partial_{x}s^{\alpha}=\partial_{x}s^{\alpha}-\frac{v}{s}\,\partial_{x}s^{\alpha}$ (this is clear when $s>0$ and when $s=0$ we have $\partial_{x}s^{\alpha}=0$ by (C6)) so that we deduce $\overline{\frac{v}{s}\,\partial_{x}s^{\alpha}}=\frac{v}{s}\,\partial_{x}s^{\alpha}$. The fact that $(u,v)$ is a weak solution to (1.2) is straightforward since $\frac{u_{\varepsilon}}{{s_{\varepsilon}}}\,\partial_{x}s_{\varepsilon}^{\alpha}$ and $\frac{v_{\varepsilon}}{{s_{\varepsilon}}}\,\partial_{x}s_{\varepsilon}^{\alpha}$ are the only nonlinear terms in (1.3). ∎
@@ 5.1. The support of $\pi_{t,x}$ is (almost always) at most a line
Note that only the last scenario (when $\operatorname{supp}\pi_{t,x}$ is contained in $\{(\lambda_{1},\mathcal{F}\,\lambda_{1})\}$ does not lead directly to (4.16) and needs further investigation in Section 5.2.
@@
Step 1. The case $s(t,x)=0$. We apply (4.2) with $f(t,x,\lambda_{1},\lambda_{2})=\mathds{1}_{s(t,x)=0}\,|\lambda_{2}|$ which is allowed since by (C6), the sequence $\{\mathds{1}_{s=0}\,|\partial_{x}s_{\varepsilon}^{\alpha}-\partial_{x}s^{\alpha}|\}$ converges strongly to 0 in $L^{1}((0,T)\times\mathbb{R})$. Hence,
@@
If $x=0$ or $y=0$, the inequality is trivial so we assume $x,y>0$. Moreover, both sides do not change when we swap $x$ and $y$ so we may assume that $x\geq y$. We let $t=\frac{x}{y}$ so dividing the inequality by $y$, it is equivalent to
@@
$\Delta:=\left|f(u_{\varepsilon},v_{\varepsilon})-f\left(\frac{u_{\varepsilon}}{s_{\varepsilon}}\,s,\frac{v_{\varepsilon}}{s_{\varepsilon}}\,s\right)\right|$
is uniformly bounded in $L^{\infty}((0,T)\times\mathbb{R})$ and $\Delta(t,x)\leq\omega_{f}(|s_{\varepsilon}-s|)\to 0$ a.e. with $\omega_{f}$ being a modulus of continuity of $f$ on some bounded set $0,M]^{2}$ such that $s_{\varepsilon}\leq M$ on $0,T]\times\mathbb{R}$. By the dominated convergence $\Delta\to 0$ strongly in $L^{2}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ so we can replace $f(u_{\varepsilon},v_{\varepsilon})$ with $f\left(\frac{u_{\varepsilon}}{s_{\varepsilon}}\,s,\frac{v_{\varepsilon}}{s_{\varepsilon}}\,s\right)$ in all the sequences whose limits appear in (5.7) (this uses that $\{\partial_{x}s_{\varepsilon}^{\alpha}\}$ and $\{u_{\varepsilon}\}$ are bounded in $L^{2}((0,T)\times\mathbb{R})$ by (E3) and (E4)). We write
@@
where $\overline{u\,\partial_{x}s^{\alpha}}$, $\overline{E}$, $\overline{E\,\partial_{x}s^{\alpha}}$, $\overline{u\,v\log v}$, $\overline{v\log v}$ are weak limits as in (C9), (C11), (C12). We apply Theorem 4.3 to the PDE for $u_{\varepsilon}$ and to the PDE for $E_{\varepsilon}=u_{\varepsilon}\log u_{\varepsilon}+v_{\varepsilon}\log v_{\varepsilon}$ given by Proposition 3.6, namely we let
$U_{\varepsilon}=u_{\varepsilon},$ $F_{\varepsilon}=u_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})+u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}],$
$V_{\varepsilon}=E_{\varepsilon},$ $G_{\varepsilon}=(E_{\varepsilon}+s_{\varepsilon})\,\partial_{x}p(s_{\varepsilon})+u_{\varepsilon}\log u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]+v_{\varepsilon}\log v_{\varepsilon}\,\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}],$
$R_{\varepsilon}=\varepsilon\,\partial^{2}_{x}u_{\varepsilon},$ $S^{1}_{\varepsilon}=h_{\varepsilon},\quad S^{2}_{\varepsilon}=g_{\varepsilon},$
where $g_{\varepsilon}$, $h_{\varepsilon}$ are defined in Proposition 3.6. We can apply Theorem 4.3 because
  • $\{U_{\varepsilon}\}$, $\{V_{\varepsilon}\}$ are bounded in $L^{\infty}(0,T;L^{2}(\mathbb{R}))$ by (E4) and (E5),
  • $\{F_{\varepsilon}\}$ is bounded in $L^{2}((0,T)\times\mathbb{R})$ thanks to (E3), (E4) (since $|u_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})|\leq\frac{1}{\alpha}\,|\partial_{x}s_{\varepsilon}^{\alpha}|$) and Assumption 1.1 which implies that $\{\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\}$, $\{\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]\}$ are bounded in $L^{\infty}((0,T)\times\mathbb{R})$,
  • $\{G_{\varepsilon}\}$ is bounded in $L^{2}((0,T)\times\mathbb{R})$ because $\{E_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})\}$ is bounded by the arguments as in (3.16) while for the other terms we argue as for $\{F_{\varepsilon}\}$ together with (E5),
  • $\{R_{\varepsilon}\}$ is strongly compact in $L^{2}(0,T;H^{-1}(\mathbb{R}))$ since $R_{\varepsilon}=\partial_{x}(\varepsilon\,\partial_{x}u_{\varepsilon})$ with $\varepsilon\,\partial_{x}u_{\varepsilon}\!\to\!0$ strongly in $L^{2}((0,T)\times\mathbb{R})$ by (E8),
  • $\{S^{1}_{\varepsilon}\}$ is strongly compact in $L^{2}(0,T;H^{-1}(\mathbb{R}))$ by Proposition 3.6,
  • $\{S^{2}_{\varepsilon}\}$ is bounded in $\mathcal{M}_{\operatorname{loc}}([0,T]\times\mathbb{R})$ by Proposition 3.6.
By Theorem 4.3, the weak limit of $F_{\varepsilon}\,V_{\varepsilon}-U_{\varepsilon}\,G_{\varepsilon}$ is $\overline{F}\,\overline{V}-\overline{U}\,\overline{V}$. We compute
$F_{\varepsilon}\,V_{\varepsilon}-U_{\varepsilon}\,G_{\varepsilon}=\,$ $(u_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})+u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}])\,E_{\varepsilon}$
$-u_{\varepsilon}\,((E_{\varepsilon}+s_{\varepsilon})\,\partial_{x}p(s_{\varepsilon})+u_{\varepsilon}\log u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]+v_{\varepsilon}\log v_{\varepsilon}\,\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])$
$=\,$ $u_{\varepsilon}(E_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\!-\!s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})\!-\!u_{\varepsilon}\log u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\!-\!v_{\varepsilon}\log v_{\varepsilon}\,\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])$
$=\,$ $u_{\varepsilon}\,v_{\varepsilon}\log v_{\varepsilon}\,\partial_{x}(\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])-u_{\varepsilon}\,s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon}).$
@@
$F_{\varepsilon}\,V_{\varepsilon}-U_{\varepsilon}\,G_{\varepsilon}\rightharpoonup\overline{u\,v\log v}\,\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v])-s\,\overline{u\,\partial_{x}p(s)}.$
The convergences above together with (C2), (C12) allow us to identify also products of the weak limits as
$\overline{F}\,\overline{V}=(\overline{u\,\partial_{x}p(s)}+u\,\partial_{x}\mathcal{V}^{1}[u,v])\,\overline{E},$
$\overline{U}\,\overline{G}=u\,(\overline{E\,\partial_{x}p(s)}+s\,\partial_{x}p(s)+\overline{u\log u}\,\partial_{x}\mathcal{V}^{1}[u,v]+\overline{v\log v}\,\partial_{x}\mathcal{V}^{2}[u,v]).$
@@
$\overline{F}\,\overline{V}-\overline{U}\,\overline{G}=\overline{u\,\partial_{x}p(s)}\,\overline{E}-u\,\overline{E\,\partial_{x}p(s)}-u\,s\,\partial_{x}p(s)+u\,\overline{v\log v}\,\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v]).$
@@
$\overline{u\,\partial_{x}p(s)}\,\overline{E}-u\,\overline{E\,\partial_{x}p(s)}+s\,\overline{u\,\partial_{x}p(s)}-u\,s\,\partial_{x}p(s)\,+\\ +(u\,\overline{v\log v}-\overline{u\,v\log v})\,\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v])=0.$
@@
so, since $\mathcal{F}\neq 0$, we obtain (5.9) as desired.
@@
see [39, Chapter 6]. We can apply (5.11) with $f(\lambda)=\lambda\log(\lambda)$ and $f(\lambda)=\lambda^{2}$ by (E4) and (E5). Now, since $\mu_{t,x}$ is a probability measure and $0,\infty)\mapsto\lambda\log\lambda$ is convex, by Jensen’s inequality
@@
$(v_{\varepsilon}-v)^{2}=(u_{\varepsilon}-u)^{2}+(s_{\varepsilon}-s)\,(v_{\varepsilon}-v+u-u_{\varepsilon})$
@@
We collect all terms in $Y^{2,1}_{\varepsilon}$ and $Y^{2,2}_{\varepsilon}$ converging at least weakly to 0
@@
Combining with $Y^{1}_{\varepsilon}=(E+s)\,(u_{\varepsilon}^{2}-u^{2})$, we obtain
$Y_{\varepsilon}=Y^{1}_{\varepsilon}+Y^{2}_{\varepsilon}\geq Z_{\varepsilon}+\frac{s-s_{\varepsilon}}{2}(u_{\varepsilon}-u)^{2}.$
We define $\widetilde{Z_{\varepsilon}}:=Z_{\varepsilon}+\frac{s-s_{\varepsilon}}{2}(u_{\varepsilon}-u)^{2}$ and we observe that $\widetilde{Z_{\varepsilon}}\rightharpoonup 0$ weakly in $L^{1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ (because it contains either terms composed of $(s_{\varepsilon}-s)$ which converge strongly by (C1) or terms composed of $(u_{\varepsilon}-u)$, $(v_{\varepsilon}-v)$ converging weakly by (C2), multiplied by some functions which are bounded independently of $\varepsilon$). Let $\varphi\in C_{c}^{\infty}((0,T)\times\mathbb{R})$, $\varphi\geq 0$. By $Y_{\varepsilon}-\widetilde{Z_{\varepsilon}}\geq 0$, we have
@@
We pass to the proof of modes of convergences. Clearly, the weak/weak convergences of $\{u_{\varepsilon}\}$, $\{v_{\varepsilon}\}$ and strong convergence of $\{s_{\varepsilon}\}$ follow directly from Proposition 3.4. We now prove the strong convergence of $\{u_{\varepsilon}\}$, $\{v_{\varepsilon}\}$ in $L^{p}(\mathcal{D})$ and a.e. on
$\mathcal{D}:=\{(t,x)\in~[0,T]~\times~\mathbb{R}~:s(t,x)=0\mbox{ or }s(t,x)>0,\mathcal{V}^{1}[u,v]\neq\mathcal{V}^{2}[u,v]\}.$
Let $\{\mu_{t,x}\}_{t,x}$ be the Young measure of $\{u_{\varepsilon}\}$, i.e. whenever $f(t,x,\lambda):[0,T]\times\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ is a Carathéodory function (measurable in $(t,x)$ and continuous in $\lambda$) and $\{f(t,x,u_{\varepsilon})\}$ is weakly compact in $L^{1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$, $f(t,x,u_{\varepsilon})\rightharpoonup\int_{[0,\infty)}f(t,x,\lambda)\mathop{}\!\mathrm{d}\mu_{t,x}\mbox{ weakly in }L^{1}_{\operatorname{loc}}((0,T)\times\mathbb{R}),$ see [39, Chapter 6]. We will prove that for $(t,x)\in\mathcal{D}$ we have $\mu_{t,x}=\delta_{u(t,x)}$ which implies the claim since $|u_{\varepsilon}-u|^{p}\rightharpoonup\int_{[0,\infty)}|\lambda-u(t,x)|^{p}\mathop{}\!\mathrm{d}\mu_{t,x}(\lambda)=0$ for $p\geq 1$. We consider two cases.
  • When $s(t,x)>0,\mathcal{V}^{1}[u,v]\neq\mathcal{V}^{2}[u,v]$, by Propositions 5.1 and 5.2, $\pi_{t,x}((\mathbb{R}\setminus\{0\})\times\mathbb{R})=0$ so that $\lambda_{1}=0$ for a.e.-$\pi_{t,x}$ $(\lambda_{1},\lambda_{2})$. In particular, by (C13) and $\overline{\,\frac{u^{2}}{s^{2}}\,}-\frac{u^{2}}{s^{2}}=\int_{\mathbb{R}^{2}}\lambda_{1}^{2}\mathop{}\!\mathrm{d}\pi_{t,x}(\lambda_{1},\lambda_{2})$,
We now discuss the strong convergence of $\{\partial_{x}s_{\varepsilon}^{\alpha}\}$. This argument is nowadays standard and has been used e.g. in [19, 43, 24]. Multiplying the PDE for $s_{\varepsilon}$ by $s_{\varepsilon}^{\alpha}$ we have for each $\varepsilon>0$ and $t\in[0,T]$
@@
We also have this identity for $s$ because we can use $s^{\alpha}$ as a test function in Definition 1.2 as $\partial_{x}s^{\alpha}\in L^{2}((0,T)\times\mathbb{R})$. More precisely, we extend $s(t,x)=s^{0}(x)$ for $t<0$, we test the PDE with $((s\ast\eta_{\delta}\ast\omega_{\kappa}\,\psi)^{\alpha}\,\psi)\ast\eta_{\delta}\ast\omega_{\kappa}$, where $\psi$ is smooth and compactly supported, $\{\eta_{\delta}\}_{\delta\in(0,1)}$ is a mollification in time while $\{\omega_{\kappa}\}_{\kappa\in(0,1)}$ is a mollification in space. After summing up and obtaining $\frac{1}{\alpha+1}\partial_{t}\int_{\mathbb{R}}(s\ast\eta_{\delta}\ast\omega_{\kappa}\,\psi)^{\alpha+1}\mathop{}\!\mathrm{d}x$, one integrates in time and we can pass to the limit $\delta,\kappa\to 0$ and $\psi\to 1$ on $\mathbb{R}$. We obtain

### 16 Apr 2026 v2

@@ 4.3. Existence of solutions in terms of $\pi_{t,x}$
Let $\pi_{t,x}$ be the Young measure of the sequence (4.1). Let $u$, $s$, $\overline{\frac{u}{s}\,\partial_{x}s^{\alpha}}$ and $\overline{\frac{v}{s}\,\partial_{x}s^{\alpha}}$ be limits as in (C1), (C2) and (C14). Then, for a.e. $(t,x)\in[0,\infty)\times\mathbb{R}$
@@
In particular, suppose that for a.e. $(t,x)\in[0,\infty)\times\mathbb{R}$
@@
This proves (4.4). Next, if (4.5) is satisfied then by (4.4), $\overline{\frac{u}{s}\,\partial_{x}s^{\alpha}}=\frac{u}{s}\,\partial_{x}s^{\alpha}$. By (C14), $\overline{\frac{v}{s}\,\partial_{x}s^{\alpha}}=\partial_{x}s^{\alpha}-\overline{\frac{u}{s}\,\partial_{x}s^{\alpha}}$ and we also have $\frac{v}{s}\,\partial_{x}s^{\alpha}=\partial_{x}s^{\alpha}-\frac{u}{s}\,\partial_{x}s^{\alpha}$ (this is clear when $s>0$ and when $s=0$ we have $\partial_{x}s^{\alpha}=0$ by (C6)) so that we deduce $\overline{\frac{v}{s}\,\partial_{x}s^{\alpha}}=\frac{v}{s}\,\partial_{x}s^{\alpha}$. Now, the fact that $(u,v)$ is a weak solution to (1.3) is straightforward since $\frac{u_{\varepsilon}}{{s_{\varepsilon}}}\,\partial_{x}s_{\varepsilon}^{\alpha}$ and $\frac{v_{\varepsilon}}{{s_{\varepsilon}}}\,\partial_{x}s_{\varepsilon}^{\alpha}$ are the only nonlinear terms in (1.4). ∎
@@ 5.1. The support of $\pi_{t,x}$ is (almost always) at most a line
Note that only the last scenario (when $\operatorname{supp}\pi_{t,x}$ is contained in $\{(\lambda_{1},\mathcal{F}\,\lambda_{1})\}$) does not lead directly to (4.5) and needs further investigation in Section 5.2.
@@
Step 1. The case $s(t,x)=0$. We apply (4.2) with $f(t,x,\lambda_{1},\lambda_{2})=\mathds{1}_{s(t,x)=0}\,|\lambda_{2}|$ which is allowed since by (C6), $\mathds{1}_{s=0}\,|\partial_{x}s_{\varepsilon}^{\alpha}-\partial_{x}s^{\alpha}|\to 0$ strongly in $L^{1}((0,T)~\times~\mathbb{R})$. Hence,
@@
If $x=0$ or $y=0$, the inequality is trivial so we assume $x,y>0$. Moreover, both sides do not change when we swap $x$ and $y$ so we may assume that $x\geq y$. We let $t=\frac{x}{y}$ so dividing the inequality by $y^{3}$, it is equivalent to
@@
$\Delta(t,x):=\left|f(u_{\varepsilon},v_{\varepsilon})-f\left(\frac{u_{\varepsilon}}{s_{\varepsilon}}\,s,\frac{v_{\varepsilon}}{s_{\varepsilon}}\,s\right)\right|$
is uniformly bounded in $L^{\infty}((0,T)\times\mathbb{R})$ and $\Delta(t,x)\leq\omega_{f}(|s_{\varepsilon}-s|)\to 0$ a.e. with $\omega_{f}$ being a modulus of continuity of $f$ on some bounded set $0,M]^{2}$ such that $s_{\varepsilon}\leq M$ on $0,T]\times\mathbb{R}$. By the dominated convergence, $\Delta\to 0$ strongly in $L^{2}_{\operatorname{loc}}((0,T)\times\mathbb{R})$, so we can replace $f(u_{\varepsilon},v_{\varepsilon})$ with $f\left(\frac{u_{\varepsilon}}{s_{\varepsilon}}\,s,\frac{v_{\varepsilon}}{s_{\varepsilon}}\,s\right)$ in all the sequences whose limits appear in (5.7) (this uses that $\{\partial_{x}s_{\varepsilon}^{\alpha}\}$ and $\{u_{\varepsilon}\}$ are bounded in $L^{2}((0,T)\times\mathbb{R})$ by (E3) and (E4)). We write
@@
where $\overline{u\,\partial_{x}s^{\alpha}}$, $\overline{E}$, $\overline{E\,\partial_{x}s^{\alpha}}$, $\overline{u\,v\log v}$, $\overline{v\log v}$ are weak limits as in (C9), (C11), (C12). We apply Theorem 4.3 to the vector fields
${\bf a_{\varepsilon}}$ $=\left(u_{\varepsilon},-u_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})-u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\right),$
${\bf b_{\varepsilon}}$ $=\left((E_{\varepsilon}+s_{\varepsilon})\,\partial_{x}p(s_{\varepsilon})+u_{\varepsilon}\log u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]+v_{\varepsilon}\log v_{\varepsilon}\,\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}],E_{\varepsilon}\right).$
Let us explain why the assumptions of Theorem 4.3 are satisfied. Clearly, both $\{\bf a_{\varepsilon}\}$ and $\{\bf b_{\varepsilon}\}$ converge weakly in $L^{2}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ to some $\bf a$} and $\bf b$} by (C1), (C2), (C5), (C9), (C11), and (C12) in Proposition 3.4. Moreover, $\operatorname{div}{\bf a_{\varepsilon}}=\varepsilon\,\partial^{2}_{x}u_{\varepsilon}$ converges strongly to 0 in $H^{-1}((0,T)\times\mathbb{R})$ because, by (E8), $\varepsilon\,\partial_{x}u_{\varepsilon},\varepsilon\,\partial_{x}v_{\varepsilon}\to 0$ strongly in $L^{2}((0,T)\times\mathbb{R})$. Regarding $\operatorname{curl}{\bf b_{\varepsilon}}$, we have $\operatorname{curl}{\bf b_{\varepsilon}}=-g_{\varepsilon}-h_{\varepsilon}$, where $\{g_{\varepsilon}\}$ and $\{h_{\varepsilon}\}$ are defined in Proposition 3.6. We know that $\{g_{\varepsilon}\}$ is bounded in $L^{1}((0,T)\times\mathbb{R})$, which compactly embeds into $W^{-1,p}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ for $p<2$ [29, Theorem 6, p. 7], while $\{h_{\varepsilon}\}$ is strongly compact in $H^{-1}((0,T)\times\mathbb{R})$. Combining these two facts, $\{\operatorname{curl}{\bf b_{\varepsilon}}\}$ is strongly compact in $(W^{1,\infty}_{0}(\Omega))^{*}$ for each bounded $\Omega\subset[0,T]\times\mathbb{R}$. Finally, the product $\{{\bf a_{\varepsilon}}\cdot{\bf b_{\varepsilon}}\}$ is uniformly integrable in $L^{1}(\Omega)$ for $\Omega$ as above. Indeed, writing ${\bf a_{\varepsilon}}=(a^{1}_{\varepsilon},a^{2}_{\varepsilon})$ and ${\bf b_{\varepsilon}}=(b^{1}_{\varepsilon},b^{2}_{\varepsilon})$, $\{a^{1}_{\varepsilon}\}$ and $\{b^{2}_{\varepsilon}\}$ are bounded in $L^{\infty}((0,T)\times\mathbb{R})$ by (E4) and (E5), while $\{a^{2}_{\varepsilon}\}$ and $\{b^{1}_{\varepsilon}\}$ are bounded at least in $L^{2}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ by (C11), (C12), and (E5). It follows that $\{{\bf a_{\varepsilon}}\cdot{\bf b_{\varepsilon}}\}$ is bounded in $L^{2}(\Omega)$, which implies uniform integrability in $L^{1}(\Omega)$.
We compute
${\bf a_{\varepsilon}}\cdot{\bf b_{\varepsilon}}=\,$ $u_{\varepsilon}\,((E_{\varepsilon}+s_{\varepsilon})\,\partial_{x}p(s_{\varepsilon})+u_{\varepsilon}\log u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]+v_{\varepsilon}\log v_{\varepsilon}\,\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])$
$-(u_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})+u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}])\,E_{\varepsilon}$
$=\,$ $-u_{\varepsilon}(E_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\!-\!s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})\!-\!u_{\varepsilon}\log u_{\varepsilon}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]\!-\!v_{\varepsilon}\log v_{\varepsilon}\,\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])$
$=\,$ $-u_{\varepsilon}\,v_{\varepsilon}\log v_{\varepsilon}\,\partial_{x}(\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])+u_{\varepsilon}\,s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon}).$
@@
${\bf a_{\varepsilon}}\cdot{\bf b_{\varepsilon}}\rightharpoonup-\,\overline{u\,v\log v}\,\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v])+s\,\overline{u\,\partial_{x}p(s)}.$
The convergences above together with (C2), (C12) allow us to identify also products of the weak limits ${\bf a}=(a^{1},a^{2})$, ${\bf b}=(b^{1},b^{2})$ as
$a^{1}\,b^{1}=u\,(\overline{E\,\partial_{x}p(s)}+s\,\partial_{x}p(s)+\overline{u\log u}\,\partial_{x}\mathcal{V}^{1}[u,v]+\overline{v\log v}\,\partial_{x}\mathcal{V}^{2}[u,v]),$
${a^{2}}\,b^{2}=-(\overline{u\,\partial_{x}p(s)}+u\,\partial_{x}\mathcal{V}^{1}[u,v])\,\overline{E}.$
@@
${\bf a}\cdot{\bf b}=-\overline{u\,\partial_{x}p(s)}\,\overline{E}+u\,\overline{E\,\partial_{x}p(s)}+u\,s\,\partial_{x}p(s)-u\,\overline{v\log v}\,\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v]).$
@@
$\overline{u\,\partial_{x}p(s)}\,\overline{E}-u\,\overline{E\,\partial_{x}p(s)}+s\,\overline{u\,\partial_{x}p(s)}-u\,s\,\partial_{x}p(s)\\ +(u\,\overline{v\log v}-\overline{u\,v\log v})\,\partial_{x}(\mathcal{V}^{1}[u,v]-\mathcal{V}^{2}[u,v])=0.$
@@
Since $\mathcal{F}\neq 0$, we obtain (5.9) as desired.
@@
see [46, Chapter 6]. We can apply (5.11) with $f(\lambda)=\lambda\log(\lambda)$ and $f(\lambda)=\lambda^{2}$ by (E4) and (E5). Now, since $\mu_{t,x}$ is a probability measure and $0,\infty)\mapsto\lambda\log\lambda$ is convex, by Jensen’s inequality
@@
$(v_{\varepsilon}-v)^{2}=(u_{\varepsilon}-u)^{2}+(s_{\varepsilon}-s)\,(v_{\varepsilon}-v+u-u_{\varepsilon}),$
@@
We collect all terms in the estimates for $Y^{2,1}_{\varepsilon}$ and $Y^{2,2}_{\varepsilon}$ converging at least weakly to 0
@@
Combining with $Y^{1}_{\varepsilon}=(E+s)\,(u_{\varepsilon}^{2}-u^{2})=(E+s)\,(u_{\varepsilon}-u)^{2}+2\,u\,(E+s)\,(u_{\varepsilon}-u)$, we obtain
$Y_{\varepsilon}=Y^{1}_{\varepsilon}+Y^{2}_{\varepsilon}\geq Z_{\varepsilon}+\frac{s-s_{\varepsilon}}{2}(u_{\varepsilon}-u)^{2}+2\,u\,(E+s)\,(u_{\varepsilon}-u).$
We define $\widetilde{Z_{\varepsilon}}:=Z_{\varepsilon}+\frac{s-s_{\varepsilon}}{2}(u_{\varepsilon}-u)^{2}+2\,u\,(E+s)\,(u_{\varepsilon}-u)$ and we observe that $\widetilde{Z_{\varepsilon}}\rightharpoonup 0$ weakly in $L^{1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ (because it contains either terms composed of $(s_{\varepsilon}-s)$ which converge strongly by (C1) or terms composed of $(u_{\varepsilon}-u)$, $(v_{\varepsilon}-v)$ converging weakly by (C2), multiplied by some functions which are bounded independently of $\varepsilon$). Let $\varphi\in C_{c}^{\infty}((0,T)\times\mathbb{R})$, $\varphi\geq 0$. By $Y_{\varepsilon}-\widetilde{Z_{\varepsilon}}\geq 0$, we have
@@
We pass to the proof of modes of convergences. Clearly, the weak/weak convergences of $\{u_{\varepsilon}\}$, $\{v_{\varepsilon}\}$ and strong convergence of $\{s_{\varepsilon}\}$ follow directly from Proposition 3.4. We now discuss the strong convergence of $\{\partial_{x}s_{\varepsilon}^{\alpha}\}$. This argument is nowadays standard and has been used e.g. in [23, 50, 28]. Multiplying the PDE for $s_{\varepsilon}$ by $s_{\varepsilon}^{\alpha}$ we have for each $\varepsilon>0$ and $t\in[0,T]$
@@
We also have this identity for $s$ because we can use $s^{\alpha}$ as a test function in Definition 1.2 as $\partial_{x}s^{\alpha}\in L^{2}((0,T)\times\mathbb{R})$. More precisely, we extend $s(t,x)=s^{0}(x)$ for $t<0$, we test the PDE with $((s\ast\eta_{\delta}\ast\omega_{\kappa}\,\psi)^{\alpha}\,\psi)\ast\eta_{\delta}\ast\omega_{\kappa}$, where $\psi$ is smooth and compactly supported, $\{\eta_{\delta}\}_{\delta\in(0,1)}$ is a mollification in time while $\{\omega_{\kappa}\}_{\kappa\in(0,1)}$ is a mollification in space. After summing up and obtaining an identity for $\frac{1}{\alpha+1}\partial_{t}\int_{\mathbb{R}}(s\ast\eta_{\delta}\ast\omega_{\kappa}\,\psi)^{\alpha+1}\mathop{}\!\mathrm{d}x$, we integrate in time and we can pass to the limit $\delta,\kappa\to 0$ and $\psi\to 1$ on $\mathbb{R}$ using the estimates on $\partial_{x}s^{\alpha}$ and $s|x|$. We obtain